# Fast dynamical similarity analysis

### v1

@@ 3 Discussion
There are also numerous avenues that can further expand the applicability of the fastDSA. Similar to the original DSA [ostrow2023beyond], a key limitation of fastDSA methods are their reliance on accurately determining embedding parameters, specifically the optimal time-delay and embedding dimension (number of delays), for constructing Hankel matrices ($H$, Eq. 2). These parameters can influence the estimation of the underlying dynamics [however, it should be noted that the choice of such parameters may not always be crucial; see Supplementary Information 14, Figure 8 of ostrow2023beyond]. In our analysis of two low-dimensional nonlinear systems (see section Sensitivity to fine changes in governing dynamics), we employed the Bayesian Information Criterion [BIC, miao2007condition] over a wide range of methods to select suitable embedding parameters (see section Optimizing delay–embedding parameters, Supp. Figure 6, and Supp. Table LABEL:Table:OptimizationDelay). However, while the BIC provides a statistically grounded approach, it is not computationally efficient, which makes the development and validation of reliable, fully automated, and computationally efficient algorithms essential. Future studies should focus on devising robust (yet fast) statistical methods for adaptive parameter selection, ensuring faithful dynamic reconstruction across diverse datasets and contexts. The same also applies to the optimization procedure, for instance, the regularization weight that we introduced earlier $\lambda$, as well as learning rate parameters used in Rim- and Land-fastDSA. Lastly, even though SVHT shows promising results, it has shortcomings in very low-noise (high-SNR) settings. SVHT chooses its cutoff from the data’s singular-value spectrum (e.g., $\tau=\omega(\beta)\!\cdot\!{\rm median}$ under unknown-noise). When noise is negligible, the “noise bulk” nearly vanishes, the median becomes signal-dominated, and the threshold drops too low. As a result, many tiny but real components—introduced by delay/Hankel embedding and our non-orthogonal geometric deformations, clear the cutoff and are counted as extra signals, inflating the estimated rank. When noise increases, the tail becomes noise-dominated again, the median rises, and SVHT returns a smaller, more reasonable rank. These effects warrant further investigation to design thresholds and preprocessing that remain stable irrespective of SNR level. Thus, further investment on tuning the hyperparameter of dynamics similarity analysis methods (beyond rank), hold great potential to widen the applicability of methods for dynamic similarity estimation.
Despite these limitations, the considerable computational efficiency of fastDSA facilitates its application in extensive, large-scale, and statistical analyses. Comparison of dynamics across multiple systems significantly benefits use cases that involve large-scale comparisons or parameter sweeps. For instance, recent work demonstrates that DSA can reveal how computations develop in recurrent neural networks by linking evolving dynamical motifs to learned task structure and performance [guilhot2024dynamical]. It highlights how architectural features and training conditions influence internal dynamics, providing a powerful lens on neural computation. With its improved efficiency, fastDSA enables such analyses at scale, supporting broader studies across architectures, input regimes, and learning curricula. Furthermore, application of dynamic similarity analysis has also recently been extended to investigate information processing in biological neural networks based on dense multielectrode neural recordings [codol2024brain, huang2025inputdsa], where large-scale statistical analysis is of paramount importance. An avenue for further extending fastDSA involves integrating Dynamic Mode Decomposition (DMD) variants tailored for noisy or real-world data. Noise-aware methods such as optimized DMD or Bayesian DMD [takeishi2017bayesian] could improve the robustness of fastDSAs when applied to empirical datasets characterized by significant noise levels. Exploring these enhancements would not only bolster the applicability of fastDSA in neuroscience and artificial intelligence but also extend its utility to a wider range of scientific and engineering fields where dynamic similarity assessment is critical. To sum, fastDSA provides a computationally efficient framework for dynamical similarity analysis, is well-suited for large-scale comparisons of neural networks, neuronal circuits, or data with a model, and its design supports future extensions to broader tasks and data modalities.
@@ 4.1.1 FastDSA with regularized optimization (RO-fastDSA)
Following the mathematical foundation of DSA [see, section Accelerating dynamical similarity analysis and ostrow2023beyond], in the first variant, we relax the orthogonality constraint on $C$ during optimization by introducing a penalty term that discourages deviations from orthogonality, and then apply a post–processing step to enforce exact orthogonality at the end. Concretely, we minimize the following regularized loss:
@@
penalizes departures from orthogonality, controlled by the parameter $\lambda>0$. We treat $\lambda$ as a hyperparameter, though it could be selected by more principled procedures [e.g., mackevicius2019unsupervised]. Using gradient descent, we update $C$ iteratively:
@@
To ensure that the resulting matrix lies in the special orthogonal group $SO(n)$, we correct the determinant of $C_{\text{orthogonal}}$ following lawrence2019purely: if $\det(C_{\text{orthogonal}})<0$, we flip the sign of the last column of $U$ (corresponding to the smallest singular value) and recompute $C_{\text{orthogonal}}=UV^{\top}$. This guarantees that $C_{\text{orthogonal}}$ is a proper rotation (no reflections) and belongs to $SO(n)\subset O(n)$. Overall, allowing $C$ to deviate slightly from strict orthogonality during optimization, smooths the loss landscape and typically speeds up convergence, while the final projection step ensures that the solution exactly satisfies the orthogonality constraint.
@@ 4.1.3 FastDSA with landing-based optimization (Land-fastDSA)
In contrast to the strict on-manifold scheme that was described above, the Landing algorithm [ablin2022fast] follows a manifold-aware optimization strategy that updates $C$ in directions consistent with the geometry of $O(n)$, but performs all steps in the ambient Euclidean space. This avoids per-step retractions while keeping iterates close to the orthogonal manifold throughout optimization. To compute these geometry-aware updates, we again exploit the local structure of $O(n)$. The tangent space of $O(n)$ at a point $C$ is
@@
Overall, by relaxing optimization to Euclidean space while using geometry-aware directions, the Landing-based fastDSA preserves DSA’s alignment objective but avoids costly per-step retractions, yielding substantial runtime savings with minimal impact on the accuracy of dynamical (dis)similarity estimation.
@@ 4.2 A heuristic optimality for a pair of Hankel embeddings
is the safer option, as it avoided sharp MSE degradation while preserving sensitivity to (possibly) weak but meaningful modes.
@@ 5 Code availability
The fastDSA package is available via GitHub at https://github.com/CMC-lab/fastDSA [behradcmclab].
@@ 7 Author contribution
8 Supplementary information
8.1 Derivations of optimization update rule for RO-fastDSA
We aim to minimize the Eq. 5 loss function with respect to a transformation matrix $C$, where the first term measures similarity between $A$ and $CBC^{\top}$, and the second term is an orthogonality regularization that enforces $C\in O(n)$. For the gradient of transformation loss term, let:
$\mathcal{L}_{1}(C)=\|A-CBC^{\top}\|_{F}^{2}=\text{Tr}(E^{\top}E),\quad\text{where }E=A-CBC^{\top}\,.$ (58)
To compute the gradient $\nabla_{C}\mathcal{L}_{1}$, using the Frobenius–trace identity $\|X\|_{F}^{2}=\mathrm{Tr}(X^{\top}X)$, we first differentiate:
$dE=-(dC)BC^{\top}-CB(dC)^{\top}\,.$ (59)
Thus, we get:
$d\mathcal{L}_{1}=2\text{Tr}(E^{\top}dE)=-2\text{Tr}(E^{\top}(dC)BC^{\top})-2\text{Tr}(E^{\top}CB(dC)^{\top})\,.$ (60)
Using the transpose/trace rule,
$\mathrm{Tr}(X)=\mathrm{Tr}(X^{\top})\,$,
$\mathrm{Tr}(X\,dC^{\top})=\mathrm{Tr}(X^{\top}dC)\,$,
and cyclic properties of trace,
$\mathrm{Tr}(ABC)=\mathrm{Tr}(BCA)=\mathrm{Tr}(CAB)\,$,
we have:
$d\mathcal{L}_{1}=-2\text{Tr}(BC^{\top}E^{\top}dC)-2\text{Tr}(B^{\top}C^{\top}EdC)\,$, (61)
$d\mathcal{L}_{1}=-2\text{Tr}\left(\left[BC^{\top}E^{\top}+B^{\top}C^{\top}E\right]dC\right)\,.$ (62)
Matching with:
$d\mathcal{L}_{1}=\text{Tr}\left((\nabla_{C}\mathcal{L}_{1})^{\top}dC\right)\,.$ (63)
Thus, we identify the gradient as:
$\nabla_{C}^{(1)}=(A-CBC^{\top})CB^{T}+(A-CBC^{\top})^{\top}CB\,.$ (64)
Then we proceed with the second term, which is defined as:
$\mathcal{L}_{2}(C)=\|C^{\top}C-I\|_{F}^{2}=\text{Tr}[(C^{\top}C-I)^{2}]\,.$ (65)
Let us denote:
$G=C^{\top}C-I\,.$ (66)
Then, we have:
$d\mathcal{L}_{2}=2\ \text{Tr}(G^{\top}dG)=2\ \text{Tr}[(C^{\top}C-I)^{\top}d(C^{\top}C)]\,.$ (67)
Lastly, using the product rule:
$d(C^{\top}C)=dC^{\top}C+C^{\top}dC\,$, (68)
and substituting:
$d\mathcal{L}_{2}$ $=2\ \text{Tr}[(C^{\top}C-I)(dC^{\top}C+C^{\top}dC)]$ (69)
$=2\ \text{Tr}[C(C^{\top}C-I)dC^{\top}]+2\ \text{Tr}[(C^{\top}C-I)C^{\top}dC]$ (70)
$=2\ \text{Tr}[(C^{\top}C-I)C^{\top}dC]+2\ \text{Tr}[(C^{\top}C-I)^{\top}C^{\top}dC]$ (71)
$=2\ \text{Tr}\left[\left((C^{\top}C-I)C^{\top}+(C^{\top}C-I)^{\top}C^{\top}\right)dC\right]\,$, (72)
we can derive the gradient:
$\nabla_{C}^{(2)}=2\left[C(C^{\top}C-I)+C(C^{\top}C-I)^{\top}\right]=4(CC^{\top}C-C)\,.$ (73)
Since $C^{\top}C-I$ is symmetric, the full gradient is:
$\nabla_{C}\mathcal{L}(C)=\nabla_{C}^{(1)}+\lambda\nabla_{C}^{(2)}\,.$ (74)
therefore the update rule is Eq. 37.
8.2 Derivations of optimization update rule for Land-fastDSA
We minimize the alignment objective
$f(C)\;=\;\bigl\|A-CBC^{\top}\bigr\|_{F}^{2},\qquad C\in O(n)\,$, (75)
and derive the update used by the Landing scheme. Let
$E\;=\;A-CBC^{\top}\,$, (76)
so that by the Frobenius–trace identity $\|X\|_{F}^{2}=\mathrm{Tr}(X^{\top}X)$ we can write
$f(C)\;=\;\mathrm{Tr}(E^{\top}E)\,$, (77)
and its first variation satisfies
$df\;=\;2\,\mathrm{Tr}\!\bigl(E^{\top}\,dE\bigr)\,.$ (78)
Differentiating $E$ gives
$dE\;=\;-(dC)\,B\,C^{\top}\;-\;C\,B\,(dC)^{\top}\,$, (79)
hence, using cyclicity of the trace and $\mathrm{Tr}(X\,dC^{\top})=\mathrm{Tr}(X^{\top}dC)$, we obtain
$df\;=\;-2\,\mathrm{Tr}\!\bigl(B\,C^{\top}E^{\top}\,dC\bigr)\;-\;2\,\mathrm{Tr}\!\bigl(B^{\top}C^{\top}E\,dC\bigr)\,$, (80)
which implies, by matching $df=\mathrm{Tr}\!\bigl((\nabla f(C))^{\top}dC\bigr)$, the Euclidean gradient in full generality:
$\nabla f(C)\;=\;-2\Bigl(E\,C\,B^{\top}\;+\;E^{\top}\,C\,B\Bigr)\,.$ (81)
When $A$ and $B$ are symmetric, this simplifies to
$\nabla f(C)\;=\;-4\,\bigl(A-CBC^{\top}\bigr)\,B\,C\,$, (82)
but note that the Landing formulation does not rely on this assumption. To move in directions compatible with the geometry of $O(n)$, we use the relative skew projection of the Euclidean gradient. Defining $\mathrm{sym}(M)=\tfrac{1}{2}(M+M^{\top})$ and $\mathrm{skew}(M)=\tfrac{1}{2}(M-M^{\top})$, the Riemannian (projected) gradient on $O(n)$ can be written as
$\mathrm{grad}\,f(C)\;=\;\nabla f(C)\;-\;C\,\mathrm{sym}\!\bigl(C^{\top}\nabla f(C)\bigr)\,$, (83)
which is equivalently expressed in “relative” form as
$\mathrm{grad}\,f(C)\;=\;\mathrm{skew}\!\bigl(\nabla f(C)\,C^{\top}\bigr)\,C\,.$ (84)
For clarity in the non-symmetric case, let $G=\nabla f(C)=-2(ECB^{\top}+E^{\top}CB)$. Then
$G\,C^{\top}\;=\;-2\Bigl(E\,C\,B^{\top}C^{\top}\;+\;E^{\top}\,C\,B\,C^{\top}\Bigr)\,,\qquad(G\,C^{\top})^{\top}\;=\;-2\Bigl(C\,B\,C^{\top}\,E^{\top}\;+\;C\,B^{\top}\,C^{\top}\,E\Bigr)\,$, (85)
and therefore
$\mathrm{skew}(G\,C^{\top})\;=\;C\,B\,C^{\top}\,E^{\top}\;+\;C\,B^{\top}\,C^{\top}\,E\;-\;E\,C\,B^{\top}C^{\top}\;-\;E^{\top}\,C\,B\,C^{\top}\,$, (86)
so the tangent component becomes
$\mathrm{skew}(G\,C^{\top})\,C\;=\;\Bigl(C\,B\,C^{\top}\,E^{\top}\;+\;C\,B^{\top}\,C^{\top}\,E\;-\;E\,C\,B^{\top}C^{\top}\;-\;E^{\top}\,C\,B\,C^{\top}\Bigr)\,C\,.$ (87)
The Landing vector field augments this tangent descent with a light attraction toward orthogonality. Let the orthogonality residual be $R(C)=CC^{\top}-I$ and define $h(C)=\tfrac{1}{2}\|R(C)\|_{F}^{2}$. A short calculation yields
$\nabla h(C)\;=\;2\,(CC^{\top}-I)\,C\,$, (88)
so the landing term
$\Lambda_{\mathrm{L}}(C)\;=\;\lambda_{\mathrm{L}}\,(CC^{\top}-I)\,C\,,\qquad\lambda_{\mathrm{L}}>0\,$, (89)
is proportional to $\nabla h(C)$ and therefore decreases $h(C)$ under explicit steps $C\leftarrow C-\eta\,\Lambda_{\mathrm{L}}(C)$ for sufficiently small $\eta>0$. Combining the tangent-like direction with the landing pull gives the complete field
$\Lambda(C)\;=\;\mathrm{skew}\!\bigl(\nabla f(C)\,C^{\top}\bigr)\,C\;+\;\lambda_{\mathrm{L}}\,(CC^{\top}-I)\,C\,$, (90)
and the explicit Landing update reads
$C^{(t+1)}\;=\;C^{(t)}\;-\;\eta\,\Lambda\!\bigl(C^{(t)}\bigr)\,,\qquad\eta>0\,.$ (91)
Because no retraction or projection is enforced at each iteration, even when the optimization begins on $O(n)$, a single explicit Euler step incurs a small drift of order $\mathcal{O}(\eta^{2})$. Thus, the optimization proceeds in the full Euclidean space, with its direction merely biased by the manifold geometry. The landing term $\lambda_{\mathrm{L}}(CC^{\top}-I)C$ acts as a soft potential that continuously reduces $\|CC^{\top}-I\|_{F}$ for small steps, counteracting this drift and keeping iterates near the manifold without enforcing feasibility at every step. If exact orthogonality is required after convergence, a single projection—such as extracting the orthogonal factor from a QR or polar decomposition—can be performed at negligible computational cost.
In settings where $C$ is parameterized as a product of factors, $C_{\mathrm{tot}}=C_{1}C_{2}\cdots C_{m}$, it is convenient to express gradients with respect to each factor. Let $L_{i}=C_{1}\cdots C_{i-1}$ and $R_{i}=C_{i+1}\cdots C_{m}$ so that $C_{\mathrm{tot}}=L_{i}\,C_{i}\,R_{i}$. A variation $dC_{i}$ induces $dC=L_{i}\,dC_{i}\,R_{i}$, and from $df=\mathrm{Tr}((\nabla f(C_{\mathrm{tot}}))^{\top}dC)$ we obtain
$G_{i}\;\equiv\;\nabla_{C_{i}}f\;=\;L_{i}^{\top}\,\nabla f(C_{\mathrm{tot}})\,R_{i}^{\top}\,$, (92)
with the corresponding tangent-like direction and landing pull
$\Psi_{i}(C_{i})\;=\;\mathrm{skew}\!\bigl(G_{i}\,C_{i}^{\top}\bigr)\,C_{i}\,,\qquad\Lambda_{\mathrm{L},i}(C_{i})\;=\;\lambda_{\mathrm{L}}\,(C_{i}C_{i}^{\top}-I)\,C_{i}\,$, (93)
leading to the per-factor explicit step $C_{i}\leftarrow C_{i}-\eta\bigl(\Psi_{i}(C_{i})+\Lambda_{\mathrm{L},i}(C_{i})\bigr)$. Working via the residual $E=A-C_{\mathrm{tot}}BC_{\mathrm{tot}}^{\top}$, one may expand
$G_{i}\;=\;-2\Bigl(L_{i}^{\top}\,E\,C_{\mathrm{tot}}\,B^{\top}\,R_{i}^{\top}\;+\;L_{i}^{\top}\,E^{\top}\,C_{\mathrm{tot}}\,B\,R_{i}^{\top}\Bigr)\,$, (94)
and substitute into the expression for $\Psi_{i}(C_{i})$. In the symmetric case $A^{\top}=A$ and $B^{\top}=B$, this simplifies to
$G_{i}\;=\;-4\,L_{i}^{\top}\,(A-C_{\mathrm{tot}}BC_{\mathrm{tot}}^{\top})\,C_{\mathrm{tot}}\,B\,R_{i}^{\top}\,$, (95)
which yields the corresponding simplified per-factor updates.
8.3 Optimizing delay–embedding parameters
Choosing the embedding parameters (the delay interval $\tau$ and the number of delays $\mu$) is a crucial step for identifying the underlying dynamics; however, it is a challenging parameter to tune. In principle, we can use well-established statistical criteria, such as BIC, and perform a grid search over $(\tau,\mu)$ to select the minimum, yielding embeddings that are dependable for downstream analysis. However, exhaustive grids over wide ranges of $\tau$ and $\mu$ are computationally expensive and often impractical. We evaluated several methods that have been suggested to provide reliable parameter selection, including AMI+FNN [wallot2018calculation], Cao’s method [cao1997practical], Joint Delay–Dimension Optimization [C–C, xu2022joint], and the Unified Embedding Algorithm [kramer2021unified].
We compared these selectors on a controlled testbed: synthetic Lorenz data ($\sigma=10$, $\rho=28$, $\beta=8/3$, $dt=0.01$), resampled to datasets of size $N\in\{1000,2000,4000,8000,12000\}$, and evaluated by downstream DMD reconstruction on a holdout set. For each method, we recorded wall time, peak memory, the chosen $(\tau,\mu)$, and reconstruction accuracy,
$\mathrm{NRMSE}\;=\;\frac{\lVert\widehat{X}-X\rVert_{F}}{\lVert X\rVert_{F}}\,.$ (96)
Scalability was summarized by fitting log–log slopes of time versus $N$ (see Supp. Figure 6 and the accompanying complexity Supp. Table LABEL:Table:OptimizationDelay).
Across our experiments, BIC (grid over $(\tau,\mu)$) delivered the lowest holdout NRMSE but at higher computational cost and with $\mu$ frequently pushed to the grid ceiling (e.g., at N=1000N{=}1000, BIC chose $(\tau{=}2,\mu{=}99)$ with NRMSE $\approx 1.5\times 10^{-5}$ in $\sim$6.3 s; scaling $\sim N^{0.80}$). A projection+ACF heuristic often matched near-BIC accuracy at a fraction of the time, making it a practical default when exhaustive grids are infeasible (e.g., at N=1000N{=}1000 on one channel it chose $(120,8)$ with NRMSE $\sim 1.9\times 10^{-3}$ in $\sim$0.25 s), though its selected lags sometimes sat at the scan boundary and performance varied across channels. While AMI+FNN and Cao’s method were consistently fast and stable (sub-second at N=1000N{=}1000; $\sim N^{0.7\text{--}0.9}$ scaling), they traded prediction fidelity for speed (e.g., NRMSE $\sim 0.18$$0.35$ at N=1000N{=}1000). The C–C criterion prioritized topological embedding over forecasting and exhibited the steepest scaling (about $N^{1.5}$), with much longer runtimes (e.g., $\sim$10 s at N=1000N{=}1000; $\sim$400 s at N=12000N{=}12000) and weaker DMD reconstruction (NRMSE $\sim 0.44$$0.73$). These outcomes reflect different objectives (forecast accuracy vs. speed vs. topology) and substantial sensitivity to data channel and sample size.
Taken together, these analyses suggest clear speed–accuracy trade-offs and non-uniform behavior across variables and $N$. At present, none of the procedures provides a universally reliable and scalable choice of $(\tau,\mu)$ across tasks, channels, and sample sizes. A principled, deformation-stable, and forecast-aware embedding selection strategy remains an open problem and warrants deeper investigation in future work.
9 Supplementary tables
Supplementary Table 1: Benchmark of delay–embedding selectors on synthetic Lorenz data. Columns: method = parameter–selection procedure; channel = state variable used ($x,y,z$); NN = number of time samples; time_sec = wall-clock runtime per run (seconds); mem_MB = peak resident memory (MB); $(\tau,\mu)$ = selected delay interval and number of delays; quality_NRMSE = normalized RMSE of DMD reconstruction on the holdout set (lower is better).
method channel N time_sec mem_MB (tau, mu) quality_NRMSE
BIC (actual) x,y,z 1000 6.31517 13.5474 (99, 2) 1.50297e-05
AMI+FNN x 1000 0.22394 0.347089 (15, 3) 0.347089
Cao’s x 1000 0.185297 0.328373 (15, 11) 0.328373
Unified Embedding Algorithm x 1000 0.253658 0.288704 (120, 8) 0.001916
C–C x 1000 10.0614 75.7413 (59, 2) 0.739594
AMI+FNN y 1000 0.22394 0.347089 (16, 4) 0.347089
Cao’s y 1000 0.185297 0.328373 (15, 11) 0.328373
Unified Embedding Algorithm y 1000 0.253658 0.288704 (120, 8) 0.001916
C–C y 1000 10.0614 75.7413 (49, 2) 0.44148
AMI+FNN z 1000 0.22394 0.347089 (16, 4) 0.347089
Cao’s z 1000 0.185297 0.328373 (15, 11) 0.328373
Unified Embedding Algorithm z 1000 0.253658 0.288704 (120, 8) 0.0019
C–C z 1000 10.0614 75.7413 (49, 2) 0.739594
BIC x,y,z 12000 408.386 1510.67 (99, 2) 1.50297e-05
AMI+FNN x 12000 1.01517 3.52901 (16, 4) 0.347089
Cao’s x 12000 0.809524 3.12453 (15, 11) 0.328373
Unified Embedding Algorithm x 12000 1.12243 1.27756 (14, 2) 0.00191
C–C x 12000 408.386 1510.67 (49, 2) 0.739594
AMI+FNN y 12000 1.01517 3.52901 (16, 4) 0.347089
Cao’s y 12000 0.809524 3.12453 (15, 11) 0.328373
Unified Embedding Algorithm y 12000 1.12243 1.27756 (14, 2) 0.00191
C–C y 12000 408.386 1510.67 (49, 2) 0.44148
AMI+FNN z 12000 1.01517 3.52901 (16, 4) 0.347089
Cao’s z 12000 0.809524 3.12453 (15, 11) 0.328373
Unified Embedding Algorithm z 12000 1.12243 1.27756 (14, 2) 0.78979
C–C z 12000 408.386 1510.67 (20, 5) 0.440172
10 Supplementary figures
List of sfigures 1 Algorithmic representations of DSA, fastDSA variants, and kwDSA: (a) DSA. (b) Regularized optimization (RO-fastDSA). (c) On-manifold Riemannian method (Rim-fastDSA). (d) Landing algorithm (Land-fastDSA). (e) kwDSA.
List of sfigures 2 Effect of the choice of the rank on reconstruction quality. MSE is computed between the noiseless latent trajectory and the reconstruction obtained by fitting DMD to a Hankel embedding of the noise-corrupted signal ($n_{\text{delays}}=68$, delay interval =1=1). The vertical dashed line marks the SVHT-selected rank $r_{\mathrm{SVHT}}$ estimated from the singular spectrum of the noisy Hankel matrix. The curve typically rises sharply for underestimation ($r<r_{\mathrm{SVHT}}$) and flattens for modest overestimation, suggesting that choosing a rank at or slightly above $r_{\mathrm{SVHT}}$ (e.g., the maximum across paired embeddings) is a safer default with respect to reconstruction error.
List of sfigures 3 Analysis of different deformation and transformation scenarios : Comparison of the loss (left) and difference scores (center) in deformation experiments between DSA and fastDSA variants across the optimization process (500 iterations). a: $\beta=1$. b: $\beta=4$. Comparison of the loss (left) and difference scores (center) in the transformation experiment between DSA and fastDSA variant across iterations. c: $\alpha=0.2$. d: $\alpha=0.9$. The Shaded regions in the line plots indicate ±1 standard error of the mean.
List of sfigures 4 Log-scale plot of Figure 5a: Similarity (distance) as a function of the parameter $\alpha$ estimated by kwDSA for four different signal-to-noise ratios (SNR = 855, 213, 8.3, 0.6) on a logarithmic scale. The plot shows the same data as the kwDSA curves in Figure 5a, but with a logarithmic y-axis to better visualize performance differences, especially at lower distances. Shaded regions in the line plots indicate ±1 standard error of the mean across 10 runs.
List of sfigures 5 Comparing fastDSA and kwDSA: Comparing clustering accuracy between fastDSA and kwDSA across different choices of the parameter $\epsilon$, using $k$-means clustering with $n_{\mathrm{clusters}}=2$, a single random initialization, and a fixed random seed to ensure reproducibility. Across combinations of $\epsilon$ for the two systems, the family of fastDSA methods consistently yields higher clustering accuracy than kwDSA, indicating greater sensitivity to subtle changes in the governing dynamics because fastDSA compares both eigenvalues and eigenvectors, whereas kwDSA focuses only on eigenvalue distributions.
List of sfigures 6 Scalability and reconstruction accuracy of delay–embedding selectors on synthetic Lorenz-based data. (a) Clock runtime versus sample size NN on log–log axes for five embedding–parameter selectors (discussed in section Optimizing delay–embedding parameters). BIC is evaluated jointly across all channels (single curve), whereas the other methods are applied per channel; curves are labeled as $\textit{method}[\textit{channel}]$ with $\textit{channel}\in\{x,y,z\}$. (b) Normalized reconstruction error (NRMSE) versus NN on log–log axes (lower is better), computed by fitting DMD using the embedding parameters chosen by each method and comparing the reconstruction to the original synthetic Lorenz data. Across all sample sizes, BIC achieves the best reconstruction accuracy but at substantially higher computational cost, whereas faster heuristics (e.g., projection+ACF, AMI+FNN, Cao’s method) scale more favorably but sacrifice predictive fidelity. No method exhibits uniformly strong performance across both speed and accuracy, underscoring that reliable and scalable delay–embedding selection remains an open challenge.

### v2

@@ 3 Discussion
There are also numerous avenues that can further expand the applicability of the fastDSA. A key limitation of fastDSA methods [similar to previous methods 57] is their reliance on accurately determining embedding parameters, specifically the optimal time-delay and embedding dimension (number of delays), for constructing Hankel matrices ($H$, Eq. 2). These parameters can influence the estimation of the underlying dynamics. In our analysis of two low-dimensional nonlinear systems (see section Sensitivity to fine changes in governing dynamics), we employed the Bayesian Information Criterion [BIC, 52] over a wide range of methods to select suitable embedding parameters (see section Optimizing delay–embedding parameters, Supp. Figure 6, and Supp. Table LABEL:Table:OptimizationDelay). However, while the BIC provides a statistically grounded approach, it is not computationally efficient, which makes the development and validation of reliable, fully automated, and computationally efficient algorithms essential. Future studies should focus on devising robust (yet fast) statistical methods for adaptive parameter selection, ensuring faithful dynamic reconstruction across diverse datasets and contexts. The same also applies to the optimization procedure, for instance, the regularization weight that we introduced earlier $\lambda$, as well as learning rate parameters used in Rim- and Land-fastDSA. Furthermore, although SVHT performs well across the regimes we investigated, it can slightly overestimate the effective rank, and this tendency is most pronounced in very low-noise (high-SNR) settings. SVHT selects its cutoff from the singular-value spectrum (e.g., $\tau=\omega(\beta)\!\cdot\!{\rm median}$ under unknown noise). When noise is negligible, the “noise bulk” effectively vanishes and the spectrum’s median can become signal-dominated, causing the threshold to drop. As a result, small components introduced by delay/Hankel embedding and by mild non-orthogonal geometric deformations can pass the cutoff and be counted as additional modes, inflating the estimated rank. In principle, a per-dataset sweep on all ranks and selects an “elbow/knee” in an MSE curve (see Figure 2c-e) could serve as a gold standard for benchmarking rank-selection rules. However, using such a sweep as a default analysis procedure is often computationally expensive (as it requires DMD-based reconstruction for the entire range of rank candidates), and in many realistic settings, the MSE–rank curve does not exhibit a clearly defined knee point, making such methods an ideal alternative. These observations motivate future work on rank-selection and preprocessing strategies that remain stable across SNR regimes and across embedding choices, while preserving the practical efficiency that enables large-scale dynamical similarity analyses. Lastly, for further extending fastDSA involves integrating Dynamic Mode Decomposition (DMD) variants tailored for noisy or real-world data. Noise-aware methods such as optimized DMD or Bayesian DMD [82] could improve the robustness of fastDSA when applied to empirical datasets characterized by significant noise levels. Exploring these enhancements would not only bolster the applicability of fastDSA in neuroscience and artificial intelligence but also extend its utility to a wider range of scientific and engineering fields where dynamic similarity assessment is critical. To sum, fastDSA provides a computationally efficient framework for dynamical similarity analysis, is well-suited for large-scale comparisons of neural networks, neuronal circuits, or data with a model, and its design supports future extensions to broader tasks and data modalities.
@@ 4.1.1 FastDSA with regularized optimization (RO-fastDSA)
FastDSA compares dynamics by aligning the induced operators/flow fields under (near-)orthogonal similarity transforms. In the first variant, we relax the orthogonality constraint on $C$ during optimization by introducing a penalty term that discourages deviations from orthogonality, and then apply a post–processing step to enforce exact orthogonality at the end. Concretely, we minimize the following regularized loss:
@@
penalizes departures from orthogonality, controlled by the parameter $\lambda>0$. We treat $\lambda$ as a hyperparameter, though it could be selected by more principled procedures [e.g., 49]. Using gradient descent, we update $C$ iteratively:
@@
To ensure that the resulting matrix lies in the special orthogonal group $SO(n)$, we correct the determinant of $C_{\text{orthogonal}}$ following Lawrence et al. [45]: if $\det(C_{\text{orthogonal}})<0$, we flip the sign of the last column of $U$ (corresponding to the smallest singular value) and recompute $C_{\text{orthogonal}}=UV^{\top}$. This guarantees that $C_{\text{orthogonal}}$ is a proper rotation (no reflections) and belongs to $SO(n)\subset O(n)$. Overall, allowing $C$ to deviate slightly from strict orthogonality during optimization, smooths the loss landscape and typically speeds up convergence, while the final projection step ensures that the solution exactly satisfies the orthogonality constraint.
@@ 4.1.3 FastDSA with landing-based optimization (Land-fastDSA)
In contrast to the strict on-manifold scheme that was described above, the Landing algorithm [1] follows a manifold-aware optimization strategy that updates $C$ in directions consistent with the geometry of $O(n)$, but performs all steps in the ambient Euclidean space. This avoids per-step retractions while keeping iterates close to the orthogonal manifold throughout optimization. To compute these geometry-aware updates, we again exploit the local structure of $O(n)$. The tangent space of $O(n)$ at a point $C$ is
@@
Overall, the Landing-based fastDSA solver retains the operator-alignment objective but replaces repeated on-manifold retractions with Euclidean-space updates guided by geometry-aware directions, followed by projection onto the orthogonal group when required. This substantially reduces computational overhead while preserving alignment quality and dynamical (dis)similarity estimates.
@@ 4.2 A heuristic optimality for a pair of Hankel embeddings
is the safer option, as it avoids sharp MSE degradation while preserving sensitivity to (possibly) weak but meaningful modes.
This analysis also supports our SVHT-based optimal rank-estimation (see section Automatic estimation of optimal rank of dynamical systems). Supp. Figure 2 suggests that slight overestimation of the optimal ranks is less harmful than rank underestimation. While modest overestimation may admit weak components (including dynamical components that correspond to noise) into the comparison pipeline, underestimation tends to discard dynamically informative modes and can produce a rapid increase in reconstruction error (MSE), degrading dynamical similarity estimates disproportionately.
@@ 5 Code availability
The fastDSA package is available via GitHub at https://github.com/CMC-lab/fastDSA [6].